OpenStudy (whovianchick):

Find (f of g)(x) when f(x)=(√x+3)/x and g(x)=(√x+3)/2x I know the answer is (f of g)(x)= (x+3)/(2x^2), I just need the steps for it, I'm doing corrections and I can't seem to do it! I know it looks hard, but if you know Pre-Calculus or know someone who does, pleeeeease help!

satellite73 (satellite73):

first off \[f(x)=\frac{\sqrt{x+3}}{x}\]?

OpenStudy (whovianchick):

yep

satellite73 (satellite73):

ok and is g the same??

OpenStudy (whovianchick):

The same except for the denominator. The denominator is 2x, rather than x

satellite73 (satellite73):

ok this is easy to explain using cut and paste

OpenStudy (whovianchick):

It's frustrating, because I know how to plug it in, the equation just seems impossible to simplify once I do. There's so many square roots.

satellite73 (satellite73):

\[(f\circ )g(x)=f(g(x))\] is always the first step

OpenStudy (whovianchick):

Yep. And you plug in the G(x) equation for all the x's in the f(x) equation

satellite73 (satellite73):

typo there \[(f\circ g)(x)=f(g(x))\]

satellite73 (satellite73):

right so this is how i do it when i type here i write \[f(g(x))=f(\frac{\sqrt{x+3}}{2x}) \]as step one

OpenStudy (whovianchick):

wait not quite. The denominator of F(x) is just x. The denominator of g(x) is 2x.

satellite73 (satellite73):

well step two actually but whatever then since \[f(\heartsuit)=\frac{\sqrt{\heartsuit}}{\heartsuit}\] i copy and past \(\frac{\sqrt{x+3}}{2x}\) in for the \(\heartsuit\)

OpenStudy (whovianchick):

ok yeah I've done that. After that I don't know what to do to simplify

satellite73 (satellite73):

oops another typo \[f(\heartsuit)=\frac{\sqrt{\heartsuit+3}}{\heartsuit}\]

satellite73 (satellite73):

ok let me copy and paste, see what we get

OpenStudy (whovianchick):

Btw thank you so much most people don't like answering advanced level questions on here!!

satellite73 (satellite73):

\[f(\heartsuit)=\frac{\sqrt{\heartsuit+3}}{\heartsuit}\] \[f(\frac{\sqrt{x+3}}{2x})=\frac{\sqrt{\frac{\sqrt{x+3}}{2x}+3}}{\frac{\sqrt{x+3}}{2x}}\]

OpenStudy (agent0smith):

@whovianchick Some people prefer to answer advanced questions on here :P

OpenStudy (whovianchick):

Yep. I've got that equation on my paper. It sucks butt.

satellite73 (satellite73):

hmm how this is going to end up as what you wrote is anyones guess in fact i think it will not, because of the nested radical is there perhaps a typo?

satellite73 (satellite73):

dividing by a fraction is the same as multiplying by the reciprocal so we can start with that

OpenStudy (whovianchick):

Here let me get a screenshot of the question

satellite73 (satellite73):

that is a good idea

OpenStudy (whovianchick):

ta-daaaaaa

satellite73 (satellite73):

lol

satellite73 (satellite73):

that is a times sign, not a circle for composition you are working too hard just multiply

OpenStudy (whovianchick):

wat

OpenStudy (whovianchick):

DANGIT

satellite73 (satellite73):

it is times, not compose do it in your head

OpenStudy (whovianchick):

IT LOOKS THE EXACT SAME IN THIS CURRICULUM

satellite73 (satellite73):

\[\circ \]\[\cdot\]

OpenStudy (whovianchick):

Lol. Well thank you. That was more help than I ever could've asked for. I'm gonna try to work it on my own, I'll tag you if I get lost again, but for both our sakes, I hope I don't lol

satellite73 (satellite73):

no prob

OpenStudy (whovianchick):

Well, I solved the case in record time. I also put on my readers.